Download
alice planning meeting n.
Skip this Video
Loading SlideShow in 5 Seconds..
ALICE planning meeting PowerPoint Presentation
Download Presentation
ALICE planning meeting

ALICE planning meeting

89 Views Download Presentation
Download Presentation

ALICE planning meeting

- - - - - - - - - - - - - - - - - - - - - - - - - - - E N D - - - - - - - - - - - - - - - - - - - - - - - - - - -
Presentation Transcript

  1. ALICE planning meeting 4.2.2011

  2. Outline • ALICE Christmas break schedule • Access UX25, L3 and MNF • Services interruption • Gas, cooling and electricity • DAQ and DCS plans • Detectors activities • Day-by-day schedule wks 49-7 • Visits • AOB ALICE planning meeting

  3. Christmas break schedule • Start on Tuesday, 7th December. End on Friday, 18th February • Plan: • Open both L3 doors • Installation of 3 TRD and 6 EMCal modules ALICE planning meeting

  4. ALICE Christmas break schedule Open L3 doors TRD Close Experiment EMCal Open Experiment Close L3 doors ALICE planning meeting

  5. Access restrictions The reasons to restrict the access are: Handling of shielding blocks Lowering of heavy structures (yellow-platform, TRD and EMCal rotators and modules) Installation of EMCal modules Removal of L3 ventilation Evacuation tests … The access restrictions concern the UX25 cavern, MNF, L3 and CRs SLIMOS shifts still missing in wks 49-50 and 4-5 Indications are given (next slides) concerning the access during the whole break – more precise info will be discussed at 5 o’clock daily meetings ALICE planning meeting

  6. UX25 access • Rules for access restrictions: • Removal/installation of shielding (restriction includes RB24 triplets) • PX24 blocks: no access CR5 and UX25 • PX24 beams (35t): no access CR1-5, Expo, and UX-25 • All other blocks: no access to UX-25 • LHC evacuation test (7/12 after 18:00 and 24/12) – no access UX25 • Gas functional test (10/12 - 2 hrs) – no access UX25 and CR1-5 • AUG test (20/1) (Arrêt Urgence General – General Emergency Stop) – no access UX25 • DSO test (12-13/1) (Department Safety Officer) – no access UX25 • No access SX2 during manipulation of blocks in SX2 (except ACR and WRs) ALICE planning meeting

  7. L3 (and MNF) access Ventil. EMCal 13 3 24 31 No Ventilation EMCal EMCal Evac. Lift No Ventilation • Wks 50 and 5: access (TPC) to be coordinated with TC (transport) • Wks 1-2: no access during EMCal SMs insertion inside L3 (see day-by-day) • When moving from low-beta to yellow platform: no access to MNF • When inserting the module: no access L3 and MNF • Watcher needed in wk 49 and from wk 4 onwards • 11/2-18/2: TPC Kr calibration: all detectors (mainly PMD, SSD, SDD, TRD and TOF) be switched on by Tue 8/2 and not change conditions anymore since then ALICE planning meeting

  8. Services interruption

  9. Maintenance gas systems New gas monitoring 13 Flowmeters Functional tests • 13-15/12: • Stop N2, CO2, Neon, Ar (24 hours max N2, CO2 and Neon) • TRD backup (N2 bottle in CR5) – HMPID backup (Ar bottle in SG2) • 12/1 (during the cooling stop): • StopC2H2F4, Isobutane, CH4 and SF6 -> will affect TOF and purifiers (no Ar/H2) ALICE planning meeting

  10. Gas system status over Christmas • TPC: flush CO2 • TRD: ON • TOF: ON – stop 11-12/12 and during cooling stop (KNF pump stopped) • HMPID: ON (Ar backup) • PHOS: N2 flushing • PMD: ON (except 11-12/12 – gas stop) • Muon TRK: backup CO2 • Muon TRG: backup Ar • No flammable gases from 7 Dec until 25 Jan PM (after maintenance UX25, Muon TRK-TRG ventilation) ALICE planning meeting Albin

  11. Maintenance cooling unitsCooling Plants and Primary Network Maintenance and consolidation L3 and Dipole TPC 10 17 24 TOF/PHOS EMC SSD/SDD 12 TPC RR Chillers Cooling towers TRD TPC HMPID 21 SPD Maintenance on production network. No primary water from 12/1 (14:00) to 21/1 (AM) CR1-2-3, the Wiener PS and detector cooling plants affected (plants can RUN, but no cooling power). DCS will be kept running at all times (dedicated chiller) Maintenance on cooling plants. Replacement of pneumatic actuators for TOF/TRD/TPC plants + TOF/PHOS flowmeter. CP will be stopped during maintenance ALICE planning meeting

  12. Maintenance ventilation units UX25, Muon TRK-TRG 17 12 SG2 L3 ALICE planning meeting

  13. Safety and electrical systems UX25 fire detect. 20 Evacuation SNIFFER mainten. PX24 fire detect. + red phones UPS UPS DSS Auto-transfert LACS doors Evacuation Evacuation Early morning/late evening (tbc) PX24 PAD 31 Lifts maintenance Guard 8-00 – 18-00 UPS Lifts maintenance AUG DSO UX25 lift ALICE planning meeting

  14. Safety and electrical systemsImpact on ALICE • LHC evacuation tests (7/12 after 18:00 and 24/12 tbc) • no access UX25 • EN/EL: interruption on auto-transfert network (9/12 from 07:00 till 07:05): • all wall outlets will be interrupted • including all non-UPS outlets in the ACR (so essentially all WRs) • including outlets SXL2, offices and barracks • non-UPS power for the detector cooling plants will be interrupted • power to all gas systems will be interrupted • the UPS powered systems will not be affected, but will run on batteries for the 2 min. interruption • the electronics racks in the cavern and CRs will not be affected. • See https://edms.cern.ch/file/1106913/1/ENNC_2010_196.pdf • DSS modification (13/1) • Will stop the full DSS, trigger all interlocks, including stopping power to all racks • EN/EL: AUG Test (20/1) (Arrêt Urgence General – General Emergency Stop) • Major electrical perturbation. Also ACR affected ALICE planning meeting

  15. DAQ upgrade plans • 21st Dec PM till the 04th Jan included • The DAQ will be OFF during the CERN closure. • The ACR PCs will be left ON to give access to the DCS for the detectors teams who need it. • 5th to the 11th of January included • DAQ restarted on the first working day for 1 week. The DAQ will be switched ON with the present version of all software. Allow a quick commissioning after the installation of new detector modules or the intervention on existing detectors. • 12th to the 23rd Jan included • The DAQ will be OFF during the cooling interruption at P2. • DAQ renovated (see TB December) • 24th January • The DAQ will be ON again. It will run under SLC5 64 bits unless there would be an unexpected key problem. • The electronic logbook will remain available during the whole period. When the DAQ will be OFF, it will be running in the DAQ lab in Meyrin. There will be a few hours of interruption to switch from P2 to Meyrin and back. ALICE planning meeting Pierre/Uli

  16. DCS plans Baseline: DCS will be operational all times where technically possible With ‘best effort’ support With access restrictions during CERN End-of-Year closure (22/12 – 4/1) Many activities and upgrades are planned Will be coordinated with individual detectors to minimise impact For central systems, as much as possible in shadow of other perturbations (e.g. cooling stop) Perturbations (e.g. for reboots etc.) cannot be excluded Detectors are invited to communicate ‘critical periods’, with services needed Will try to cater for detector needs as much as possible Known perturbations: Tue 7/12 10:00 – 12:00, interruptions of the GPN link to/from P2 Upgrade of router, test impact on ALICE Wk50 (13-17/12) – major updates to all PC (many perturbations to be expected) 12/1 – 21/1 (cooling stop), only minimal central services. All detector DCS switched off, except for special requests and if cooling power allows More detailed planning at weekly and daily meetings ALICE planning meeting André / Peter

  17. DCS plans (details) Wk 50 Mon 13/12 – Fri 17/12 (daytime): SW upgrades and patches to all systems (Operating system and PVSS) Delayed/extended due to problems with automatic PVSS upgrade procedure (several manual operations needed) Systems unavailable for short (~hrs) periods, frequent reboots Support to detectors restarting/testing their systems after upgrades Detectors shall test functionality as much as possible Wk 51 Contingency for system upgrades Support to detectors restarting/testing their systems after upgrades Detectors shall test functionality as much as possible Tue 21/12 PM: Limit access to DCS cluster during end-of-year closure Register with your SRC if you require access (see message Peter 2/12/2010) Wk 52 – Wk 01 Only ‘best effort’ support during CERN closure Wk 02 Wed 12/01 – (cooling intervention) shutdown of detector systems except special requests, within limits of cooling power ALICE planning meeting André / Peter

  18. Access requestCavern and surface • You should submit an ADI and a task procedure Attach task procedure Work breakdown structure – risk associated at each step Two persons minimum ADI https://edh.cern.ch/Document/ADI Task procedure Template EDMS 1103355 ALICE planning meeting

  19. Detectors activities

  20. EMCal Status of modules: 6 modules in SXL2 Test station installed in SXL2 Requiring a lot of access to connect everything up, test, and debug as they commission the new SMs Low voltage, high voltage (CR4), and DAQ operating as much as possible during the installation period in January. ALICE planning meeting

  21. EMCalSurveys • Zero measurement (wk 51-52), after insertion of 1st module and after the insertion of 2nd module: • A-side: CALframe feet and rails + modules (adaptors) • C-side (photogrammetry - 1/2 day) • After insertion of 3rd, after insertion of 4th and after insertion of 5th module: • A-side: points chosen depending on results previous survey • C-side: with theodolite from A-side (or full photogrammetry, depending on results previous survey) • After insertion of 6th module: • A-side: full • C-side: full ALICE planning meeting

  22. PMD • Wk 50: • Lift CROCUS (in order to open L3 doors) - can be done by P2 technicians • disconnect cables (after 16/12) • Wk 51 (20/1): move PMD in parking position (Pieter) • Wks 51-3 SPIE: new air cooling system – major installation work in January (should be done by end January) • Wks 51-4: repair electronics chains • Need Ar/CO2 at all times • Wk 4: 24/1 CROCUS platform re-installed • Wk 4 (normal hours if compatible with L3 + PMD duct installation, otherwise after 18:00): access to CROCUS • Wk 5 (after 18:00): access to CROCUS; 31/1 PM: start new cooling system; 1-2/2: test detector with new cooling; 3/2 (or later, in wk 6): move PMD to final position • Wk 7: survey of PMD ALICE planning meeting

  23. TOF/PHOS • TOF: • HV hardware checks on SM13 (December) • TRM replacement (14 TRMs on the A side) (3/01 – 12/01) • Replacement DC-DC (with no LV spare channels): • 16/12 (after 18:00) • PHOS: (A.Vinogradow at CERN 1/12 end January) • Investigate module 2 pedestal issues • Try to understand behavior of unstable branch in Module 3 • Remake ISEG crate to connect to the hardware interlock • Maintenance crystal cooling plant (January) ALICE planning meeting Gilda/Yury

  24. Muon TRG • RPC Detector/gas system: • Check gas leak on few RPC detectors and repair 2 gas distribution channels. Replacement of 1 HV cable. From Monday 10/01 (afternoon) to Friday 14/01 • Upgrade Sensor gas humidity and pressure regulator (CR5, schedule to be defined with the gas team) • Replace few RPC modules with high current: under discussion • Will restart the flammable gas flow at end of January/beginning of February, to check the detectors with HV at least 15 days before the first beam injection • Electronics: • Maintenance Local decision electronics (December 8th after 18:00) Done • Upgrade FE threshold distributors, higher segmentation (level 4 gangway, December 8th after 18:00) Done • Upgrade firmware DARC and Regional board in order to speed up MTR decision electronics configuration (wk 6, after the DAQ upgrade) ALICE planning meeting Pascal/Ermanno/Paolo/Anna

  25. Muon TRK plans • December: CROCUS firmware upgrade (protection against spurious busy reset signal): • Access to Crocus, Dipole and FASS. Need nacelle wk 49 Done • High occupancy bus patches: • Remove one "bad" slat (and replace by one spare) to test it in lab Done • Access to Dipole and to muon side with nacelle in December Done • Fix all remaining “bad” slats in January, wks 2-5. Need local DAQ setup (2 tables) + ITECO • 13 January: replacement 1 quadrant ST2 (faulty wires) Done • Test spare quadrants in lab (December) ALICE planning meeting Ermanno/Herve

  26. ZDC plans • Check platform displacement with survey • One week – depending on surveyors availability ALICE planning meeting Paolo

  27. SPD cooling plans • Wk 49 (8/12 after 18:00): cut pipes in order to open the doors • Wks 2-3 (during cooling stop): wash 4 bad sectors (machine installed on I-side, close to HMPID pump). Access to the outer side of MNF (side I) to work on the heater of sector 0 (nacelle or scaffolding needed) • Wks 3-4 (as soon as the door is closed): reconnect temporarily pipes and test with dummy sector inside MNF (in parallel with shielding in wk 4) ALICE planning meeting

  28. SSD/SDD SSD: • Ton: remove a cooling pipe from SSD corrosion test setup (week 2) • Replacement CAEN module (C-side) + repair bad temperature signal at the C-side (week 2) SDD: • Calibration of 78 pressure sensors in L3 and on miniframe (January 12 to 14) ALICE planning meeting

  29. T0 • Activities planned from 7-19 December (after 18:00): Done • Check/repair the amplifiers from A and C side (inside L3) • Check the signal cables from A-side • Check/repair fast electronics modules (racks O18-19) • Need possibility to do standalone runs in parallel ALICE planning meeting

  30. TRD installation - weeks 51, 52 • 4 modules at P2 (SM2, SM11, SM 15 and SM16) • Installation sequence: 16 → 15 → 11 (or 16 → 11 → 15) • 2 technicians (Philippe and Didier) + Survey (Nail from 6/12) + TC permanently at P2 • Connection done right after installation • Survey: • Space-frame deformations and displacement of beampipe (flange), miniframe and TPC feet. Before-after each insertion • Modules (2 points/SM on 4/7 SM) in wk 50 • Modules (3 points/SM on 10/10 SM) after EMCal (wk 3) ALICE planning meeting

  31. TRD • Week 49: • LV Wiener ~70 PSU Back-Plane modifications (to improve mechanical stability, same as TPC applied already) • Netgear switches replacements (1 day) • LV line fixation: installation of improved patch panels on front of module (C-side) and replacing cables • Repair pretrigger (A-side, below PMD) – after 20/12 • Week 1: • Finalizing connections, start commissioning • Week 2: • Stable running with cosmics • 3 to 4 weeks in January or February: • Stable running with cosmics for alignment and with Krypton for gain calibrations (to be still discussed) ALICE planning meeting

  32. TPC • Recover all water inside cooling plant (week 50) • Replacement of 15 damaged FECs (equally distributed A-C sides) • ~3 days per FEC. Therefore, we would like to use all available time slots where access to the L3 is possible • At the end of this activity we will need a period of about 2 weeks where all aforesaid systems are available for final verification, and access to the TPC in case something is not working correctly. • Kr data taking at the end of the shutdown, in wk 7 ALICE planning meeting

  33. TPCReplacement of the laser • Start the dismantling of the Spectron laser (week 50, Th-Sat after 18:00): • dismount back wall of laser hut (help of P2 technician) • stop cooling water (help of P2 technician) • prepare new water connection (electrical power should be OK) • Install new Ekspla laser (week 1): • remove old laser and install new (use crane to transport laser) • connect cooling water (help of P2 technician) • reinstall back wall (maybe after cutting a hole in it) • Testing and alignment of new laser (week 2): • possibility to run laser behind closed doors (inside hut) • maybe install laser tube temporarily and shoot into TPC whenever • compatible with other installation work • Possible further testing and adjustment (week 3) ALICE planning meeting Borge

  34. ACORDE • Check some ACORDE modules (PMTs optical conections, connectors, etc) above the magnet • The plan is to start this activities on January 21st ALICE planning meeting

  35. Other activities • SPIE + CV: refurbishment ITS dehumidifier in CR – February • Replacement part at P2 in January. Installation – transport needed • Recirculation air - SPIE • New air-flowmeter (connect to ELMB: Frederick) • EN-EL (T.Bourrel): motorize L3-Dipole-Miniframe circuit breaker. • VIC 16/12 10:30 – Intervention 11/1 whole day • SPIE: cabling UX25-CR4 – 17/1 • New magnet emergency buttons + flashing lights L3 and dipole • Frederick: humidity sensor L3 ventilation (extraction) + redo CTP connectors • Intervention on cavern lift. From 28/1 17:00 to 29/1 (whole day) • One day stop – no access to UX25 ALICE planning meeting

  36. Other activities Google propose to produce a custom Streetview application for CERN showing both outside and underground areas Visit Tuesday 21 December at 16:15 18/1 filming Philippe: replace rack filters during cooling stop. Install temporary line PHOS cooling (PS rack and crystal cooling plant). Clean all WIENER PS racks with special machine ALICE planning meeting

  37. Magnet control upgrade • Wk 49: nous devons intervenir dans la MCR pour: • changement des alimentations 24V et 48 V des systèmes MSS et MCR (préventif) • changement des modules convertisseurs PT100 des hots points du Solénoïde (correctif suite à deux incidents ayant entrainés l'arrêt du Solénoïde) et en UX25, comme Philippe Bouvier l'a déjà signalé: • recâblage des AU du dipôle et du Solénoïde • signalisation du champ magnétique du dipôle et du Solénoïde • Nous étudions toujours une solution pour le monitoring des températures du Solénoïde. • L'arrêt de fin d'année va permettre de faire une meilleure approche de ce qui est existant en UX25. • En ce qui concerne l'alimentation électrique de la MCR, nous dépendons actuellement d'une seule source de courant sous UPS avec secours diesel. • Afin d'établir une redondance pour l'alimentation de la MCR je souhaiterais installer une seconde source d'alimentation sous UPS. • Je sais qu'une étude est faite à ALICE pour augmenter la puissance des UPS existants en 2012. • Je souhaiterais que cette demande soit incluse pour le calcul de puissance des nouveaux UPS (besoin pour la MCR: 16A en monophasé) ALICE planning meeting Pascal

  38. Day-by-day schedule UX25 and MNF - Wk 49

  39. Week 49 - 6/12 MondayProgram • 18:00: end HI program • Ramp down both magnets. !!! Emergency dump !!! • 18:00 – 22:00: access system test. No access UX25 • 22:00: RPE survey (Jiri + 2nd person) • Then go to restricted mode ALICE planning meeting

  40. It H Activities in UX25 - Week 49 - 6/12 Monday No access UX-25 - possibility of access from 22:00 TRD team at CERN: 6 to 10/12 ITECO HAULOTTE ALICE planning meeting

  41. Week 49 - 7/12 TuesdayProgram • 8:00 - 18:00: UX25 cavern in Restricted mode • No access UX25 (except Transport, VAC, TC and RP) • 8:30 - 9:30: RP liquid (and gas) sampling (Adeline, Pierre) • Collect RP samples (Fernando, Philippe) • Survey L3 and Dipole • 18:00 - 21:00: evacuation test – No access UX25 • From 21:00 - Cavern in General mode • 14:00 - 15:00: Portuguese Press visit. 10 persons ALICE planning meeting

  42. Week 49 - 7/12 TuesdayActivities • ALTEAD: PX24 shielding blocks – top layer removed • 8:00 to 18:00: no access CR5 • Reinstall crane electronics - Problems (ongoing) • 14:00 to 15:00: put SM16 into the rotator - Ongoing • VAC: Neon injection preparation - Started • Transport needed (EDH 4472100) – Rack moved by hand • Remove “green fence” (P2 technician) - Done • Stop flammable gases i-C4H10 and CH4 (Ar/H2 on 10/12) - Done • 8:30: Lock L3+Dipole power converter + compensator magnet - Done • Test transmission of alarm lift to FB - Done • 8:30: inspection PHOS cooling plant (Arturo, Dominique, Philippe, Sasha) – Done (will install temporary line) • 15:15: VIC SPD – Done + Gas VIC • DCS: 10:00 – 12:00, interruptions of the GPN link to/from P2 – Done. DCS got stuck. Investigating why • MCR (Magnet Control Room): HW modifications (Pascal B.) – Started • TRD (after 21:00): Wiener PS back-plane modification and LV consolidation (L3) ALICE planning meeting

  43. It H Activities in UX25 - Week 49 - 7/12 Tuesday TRD Wiener PS PX24 shielding: blocks VAC RP samples TRD Wiener PS ALICE planning meeting

  44. Week 49 - 8/12 WednesdayProgram • 8:00 - 18:00: UX25 cavern in Restricted mode • No access UX25 (except Transport, VAC, TC and RP) • From 18:00 - Cavern in General mode • 18:00: Confined Space Training • Tour in L3 at the end of the course ALICE planning meeting

  45. Week 49 - 8/12 WednesdayActivities • ALTEAD: PX24 shielding beams (35t) - Done • 8:00 to 16:30: no access CR1-5 and Expo • Move ITECO C-side - Ongoing • EN-HE: Reinstall crane electronics – Done. Crane operational • VAC: Neon injection – injection table prepared. Neon injected tomorrow afternoon • Annular ion pump (PI.45.1L2.X) switched off Activities after 16:30: • EN-CV: empty TPC tank - Ongoing • TPC: replacement of damaged FECs (C-side) - Ongoing • TRD: pretrigger PS work done. Wiener PS back-plane modification (SXL2, almost done). Start re-installing Wieners tonight. • SPD: cut inox pipes - Done • Muon TRK: CROCUS firmware upgrade • Muon TRG: Maintenance Local decision electronics and upgrade FE threshold distributors – Done • T0: investigation broken amplifier (MNF and L4-rack) - Ongoing ALICE planning meeting

  46. It Activities in UX25 - Week 49 - 8/12 Wednesday Move TRD Wiener PS T0 PX24 shielding: beams (35t) VAC Muon TRK: CROCUS firmware upgrade TPC FEC and TRD LV Cut SPD pipes TPC cooling Muon TRG TRD Wiener PS ALICE planning meeting

  47. Activities in MNF - Week 49 - 8/12 Wednesday T0 ALICE planning meeting

  48. Week 49 - 9/12 ThursdayProgram • 8:00 - 18:00: UX25 cavern in Restricted mode • No access UX25 (except Transport, VAC and TC) • From 18:00 - Cavern in General mode • Interruption on auto-transfert network (from 07:00 till 07:05). No access underground • all wall outlets will be interrupted • including all non-UPS outlets in the ACR (so essentially all WRs) • including outlets SXL2, offices and barracks • non-UPS power for the detector cooling plants will be interrupted • power to all gas systems will be interrupted • the UPS powered systems will not be affected, but will run on batteries for the 2 min. interruption • the electronics racks in the cavern and CRs will not be affected. • See https://edms.cern.ch/file/1106913/1/ENNC_2010_196.pdf ALICE planning meeting

  49. Week 49 - 9/12 ThursdayActivities • ALTEAD: MNF shielding – 2/4 “T” blocks removed • 8:00 to 18:00: No access CR1-5 during lifting up of 5 big beams • Stabilizers (after 18:00) - OK • Move ITECO I-side • VAC: Neon injection – Beampipe under Neon • RAMSES: removal of Data Acquisition modules CR5 - Done • EMCal team at CERN: fit rollers on all SMs (SXL2). Start 13:30 - OK • Dipole-Solenoid cooling stopped (Philippe) • Purge TRD cooling flexibles + replace o-rings (Philippe + Frederick) - Done Activities after 18:00: • TPC: replacement of damaged FECs (C-side) • TRD: Wiener PS back-plane modification and LV consolidation (SXL2) • Muon TRK: CROCUS firmware upgrade • T0: investigation broken amplifier (MNF and L4-rack) ALICE planning meeting

  50. It H H Activities in UX25 - Week 49 - 9/12 Thursday TRD Wiener PS T0 Stabilizers MNF shielding VAC Muon TRK: CROCUS firmware upgrade TPC FEC and TRD LV Stabilizers TRD Wiener PS Move No access CR1-5 during lifting up of 5 big beams ALICE planning meeting